If the potential across two parallel plates, separated by 1.50 cm, is 12.0 volts, what is the electric field strength in volts per meter?
1
Expert's answer
2017-06-10T05:54:11-0400
We can find the electric field strength from the formula:
E=V/d,
here, E is the electric field strength, V is the potential across two parallel plates, d is the separation between the plates.
Then, we get:
E=V/d=(12 V)/(0.015 m)=800 V/m.
Answer:
E=800 V/m.